Homework zo P. 200 Given dy 3x2+4+2 y (6)=-| dx 2ly-1)
Answer #1

. Given dy 3x² + 4x+2 2(4-1) 7(0) = dx 0 i. (1-1) dy = (3x?+ 4x + 2) da Integrate both side, we get f 4-1) dy = f(3x?+ 4x +2]